OpenStudy (anonymous):

Verify the identity. sin 4u = 2 sin 2u cos 2u

OpenStudy (anonymous):

i have two other identities i need help with if you are good with them. im pretty much awful at identities so i really need help

OpenStudy (anonymous):

This follows directly from the double angle formula: \[\sin 2 \theta = 2 \sin \theta \cos \theta\] Plug in theta=2u and you have your answer.

OpenStudy (anonymous):

but 2 * 2sin x cos x equals 4 sin x 2 cosx not 2 sin 2u cos 2 u right? or are they the same thing

OpenStudy (anonymous):

You're plugging in 2u for theta in the identity shown. This will give you 2sin(2u)cos(2u)

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

Take this identity: \[\sin 2 \theta = 2 \sin \theta \cos \theta\] and plug in 2u for theta: \[\sin 2(2u) = 2 \sin (2u) \cos (2u)\] \[\sin 4u = 2 \sin 2u \cos 2u\]

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

sin 4u = 2 sin 2u cos 2u sin 2u sin 2u = 2 sin 2u cos 2u 2sin u cos u 2 sin u cos u = 2 sin 2u cos 2u 2 sin 2u cos 2u = 2 sin 2u cos 2u does this prove it correctly?

OpenStudy (anonymous):

If you've learnt sin(A+B) = sinAcosB + cosAsinB, sin(4u) = sin(2u+2u) = sin(2u)cos(2u) + cos(2u)sin(2u) = 2 sin(2u) cos(2u).

OpenStudy (anonymous):

see that makes much more since, thankyou